Experiment 3 Prelaboratory Assignment Empirical Formulas Date Lobo Sec. Name Desk No. 1. Elemental mercury was first discovered when a mercury oxide was decomposed with heat, forming mercury metal and oxygen gas. When a 0.204-g sample of the mercury oxide heated. 0.189 g of mercury metal remains. Note: not attempt this experiment in the laboratory because of the release of toxic mercury vapor. a. What is the mole ratio of mercury to oxygen in the sample? b. What is the empirical formula of the mercury oxide? 2. A 5.90-g sample of titanium chemically combines with oxygen gas to form 9.84 gofa titanium oxide. a. What is the empirical formula of titanium oxide? b. What is the percent by mass of titanium and the percent by mass of oxygen in the sample? 3. a. Experimental Procedure, Part A. List two reasons for using crucible tongs to handle the crucible and lid after their initial firing. b. Why is it best to cool the crucible and lid (and sample) a desiccator rather than on the laboratory bench? in 15b Experiment 3 61
media%2F0f0%2F0f0ef997-16ba-48d2-96a9-b1
Answer #1

a) mercury oxide mass = 0.204 g

mercury mass = 0.189 g

oxygen mass = 0.204 - 0.189 = 0.015

mercury moles = 0.189 / 200.59 g/mol = 0.00094 mol Hg atoms

oxygen moles = 0.015 g / 16 g/mol = 0.00094 mol o atoms

Hg : O ratio = 1: 1

b) empirical formula = HgO

2) a) 5.9/47.9 = 0.123mols of Ti

9.84-5.9 = 3.94g of O2 so

3.94/16 = 0.246 mols of O2

dividing by the smallest number

0.123 / 0.123 = 1 and 0.246 / 0.123 = 2

the empirical formula is TiO2

b) molar mass of TiO2 = 79.9g/mol

% Ti = 47.9 / 79.9*100 = 59.9 %

% O = 32 / 79.9*100 = 40.1 %

3) a) 1. to prevent burning hands

2. prevent contamination

b) so it will not burn the beach